dude:

\((\frac{2}{(2-x)}+3)(\frac{x+1}{x+1})\) Find \(\lim_{x \rightarrow (-1)^-}(f(x)) \)

dude:

@Pixel

Pixel:

uhhhhhhhhhhhhhhhh

Pixel:

my question is a bonus challenge question so i don't know that subject sorry my guy

dude:

Minor change lol \(f(x)=(\frac{2}{(2-x)}+3)(\frac{x+1}{x+1})\) Find \(\lim_{x \rightarrow (-1)^-}(f(x)) \)

Nnesha:

\[f(x)=(\frac{2}{(2-x)}+3)(\frac{x+1}{x+1})\] like this ?

Nnesha:

\[\frac{\cancel{ x+1 }}{\cancel{ x+1 }}=1\]

Nnesha:

and then direct substitution